OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#10420 — anj75 bas75 bdv75 bea75 gut75 pro75 tui75 vol75
Incident Report for Web Cloud
Resolved
A fibre is cut on the Paris optic loop.
Securisation is handling it, no impact for the customers.

We have opened a ticket with our fibre operator to restore the securisation quickly.

Update(s):

Date: 2014-03-12 15:26:47 UTC
After soldering the affected fibre, the signal is back to normale.
The loop is now secure.

Date: 2014-03-12 13:19:41 UTC
The cut was repaired 930m from pro75. The team will intervene to correct the issue this afternoon.

Date: 2014-03-12 12:04:27 UTC
The gut75 DSLAM is accessible again.

Date: 2014-03-12 12:04:23 UTC
The teams are now intervening on the gut75 fibre optics.
The DSLAM has been isolated and is unreachable.


Date: 2014-03-12 09:50:34 UTC
The pro75 DSLAM is accessible again.
The intervention continuse to restore the securisation.

Date: 2014-03-12 08:37:42 UTC
The teams are intervening on the pro75 site which is currently unreachable.
Posted Mar 12, 2014 - 04:53 UTC